Hi there and thanks in advance. Betty and Karen working together can paint a house in two-thirds time it takes Karen working alone. Betty takes 6 hours to paint a house alone. How long does it take Karen to paint a house working alone?
I know the usually way to solve these but the 2/3 part is screwing me up.

Answer by jorel1380(3719)

(2/3)(1/6+1/n)=1/n
2/18+2/3n=1/n
1/9+2/3n=1/n
n+6=9
n=3
It takes Karen 3 hours working alone to paint a house
1/6+1/3=1/6+2/6=3/6=1/2
2/3X1/2=1/3..